Matt is four years younger than twice his cousin Andy’s age. The sum of their ages is greater
than 35. What is the youngest age that Andy can be?

Answer by ankor@dixie-net.com(22740)

let m = Matt's age
let a = Andy's
:
Write an equation for each statement
:
Matt is four years younger than twice his cousin Andy’s age.
m = 2a - 4
The sum of their ages is greater than 35.
m + a > 35
What is the youngest age that Andy can be?
replace m with (2a-4) from the 1st statement
2a - 4 + a > 35
3a > 35 + 4
a > 39/3
a > 13
least integer has to be a = 14
:
:
Check, find m
m = 2(14) - 4
m = 24
24 + 14 = 38 greater than 35
If a = 13
m = 2(13) - 4
m = 22
22 + 13 = 35 which is not greater than 35